BIOL 3408 Problem: Bohr NAME SHOW ALL WORK, including show on the graph how you determined the answers. Use the correct units. What is the partial pressure of oxygen at sea level for dry air? What is the partial pressure of oxygen in the outside air if saturated with water vapor at 37 degrees C (water vapor partial pressure The alveoli in your lungs are saturated with water vapor (partial pressure In addition, carbon dioxide levels are elevated (partial pressure 45 mm Hg) and the oxygen levels therefore reduced by this amount? What is the partial pressure of oxygen in your alveoli? 1. 2. 47 mm Hg)? 3. 47 mm Hg). 100 r 90 80 e 70 60 E 50F 4아 20 10 0 20 40 80 100 mm Hg 12 kPa Partial pressure of O2 in blood Using the oxygen dissociation curve, determine the saturation % of hemoglobin in the lungs Blood leaving the lungs carries 20 ml 02/100 ml. If there is no Bohr shift, and the oxygen pressure in the tissues is 35 mm Hg, how much oxygen (in ml 02/100 ml) is unloaded to the muscles? With the Bohr shift, how much is unloaded (in ml 02/100 ml)? 4. 5.

Need help with #3 please.
Answer #1

We can answer this using the Alveolar gas equation.

p_{A}O_{2}\approx F_{I}O_{2}(P_{{ATM}}-pH_{2}O)-{\frac {p_{a}CO_{2}}{RER}}

where,

pAO2- The alveolar partial pressure of oxygen (pO2) 107 mmHg (14.2 kPa)
FIO2- The fraction of inspired gas that is oxygen (expressed as a decimal). 0.21
PATM- The prevailing atmospheric pressure 760 mmHg (101 kPa)
pH2O- The saturated vapour pressure of water at body temperature and the prevailing atmospheric pressure 47 mmHg (6.25 kPa)
paCO2- The arterial partial pressure of carbon dioxide (pCO2) 40 mmHg (5.33 kPa)
RER- The respiratory exchange ratio 0.8

by using the above formula- 0.21(760-47) - 45/0.8

= 0.21(713)- 56.25

= 149.73-56.25

=93.48

So the alveolar partial pressure of oxygen will be 93.48 mmHg.

As the CO2 concentration has been increased, the curve will shift to the right. So, the % saturation for Hemoglobin will be around 95%